OpenStudy (anonymous):

How many liters of helium are present in a balloon if the balloon contains 2.23 g of He? A) 12.5 L He B) 6.25 L HE C) 22.4 L He D) 4.0 L He

OpenStudy (joannablackwelder):

1 mol of ideal gas = 22.4 L Do you know how to convert from grams to moles, then moles to liters?

OpenStudy (aaronq):

you need to be given condition specifications (temp, pressure)

OpenStudy (joannablackwelder):

True. I was assuming that a balloon would be at standard temp and pressure.

OpenStudy (theeric):

@emily.dossen Were you given the density of the helium?

OpenStudy (theeric):

If you were, let the density be represented by the variable \(\rho\). Density is mass per volume, right? So \(\rho=\dfrac{m}{V}\).

OpenStudy (isaiah.feynman):

@theEric that's what I thought too.

OpenStudy (theeric):

Or a grams/mol number, which would be on the internet...

OpenStudy (ybarrap):

For an ideal gas, $$ PV=nRT V=\cfrac{nRT}{P} $$ http://en.wikipedia.org/wiki/Pv%3Dnrt Where n is the number of moles, T is Temp in Kelvin and P is pressure in Pascals. R is the Gas Constant: \(R=8.314\,4621~\frac{\mathrm{J}}{\mathrm{mol~K}}\) http://en.wikipedia.org/wiki/Ideal_gas_constant There are 4.002602 grams per mole for He. So 2.3g He is $$ 2.23g \times \cfrac{1~ mole}{4.00260~ g}=0.557 moles $$ Therefore, Volume is $$ V=\cfrac{0.577\times 8.314\,4621 \times T}{P} $$ At 1 atm, P = 101 325 Pa (pascals) At room temperature, 25\(^\circ C\)=298.2 K So, $$ V=\cfrac{0.557\times 8.314\,4621 \times 298.2}{101\,325}\\ \approx 0.0136~m^3 $$ 1 cubic meter is 1000 Liters. So, $$ V=0.0136\times 1000=13.6~ L $$ You'll need to adjust this for your given temperature and pressure. The molar volume of an ideal gas at 1 atmosphere of pressure is 22.414 L/mol at 0 °C 24.465 L/mol at 25 °C So, at 0 °C The volume of He is, $$ 22.414 \times 0.557 = 12.5 L $$ Based on this, it appears that your He gas is probably at 0 °C and 1 atm. So, $$ V=\cfrac{0.557\times 8.314\,4621 \times 273.2}{101\,325}\\ \approx 0.01248~m^3\\ \approx 12.5 L $$